OpenStudy (anonymous):

In a right triangle, the acute angles measure x + 15 and 2x degrees. What is the measure of the smallest angle of the triangle? 15 25 40 50

Parth (parthkohli):

One angle is 90 degrees we know for sure(right triangle) they add up to 180 degrees. \(\Large \color{MidnightBlue}{\Rightarrow 90 + x + 15 + 2x= 180 }\) \(\Large \color{MidnightBlue}{\Rightarrow 105 + 3x = 180 }\) \(\Large \color{MidnightBlue}{\Rightarrow 3x = 75 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 25 }\) One angle is x + 15 = 25 + 15 = 40 Second one is 2x = 2*25 = 50 Which is the smaller one?

OpenStudy (anonymous):

40